# Robust Fast Adaptation from Adversarially Explicit Task Distribution Generation

### v1

@@ Abstract.
Meta-learning is a practical learning paradigm to transfer skills across tasks from a few examples. Nevertheless, the existence of task distribution shifts tends to weaken meta-learners’ generalization capability, particularly when the task distribution is naively hand-crafted or based on simple priors that fail to cover typical scenarios sufficiently. Here, we consider explicitly generative modeling task distributions placed over task identifiers and propose robustifying fast adaptation from adversarial training. Our approach, which can be interpreted as a model of a Stackelberg game, not only uncovers the task structure during problem-solving from an explicit generative model but also theoretically increases the adaptation robustness in worst cases. This work has practical implications, particularly in dealing with task distribution shifts in meta-learning, and contributes to theoretical insights in the field. Our method demonstrates its robustness in the presence of task subpopulation shifts and improved performance over SOTA baselines in extensive experiments. The project is available at (https://sites.google.com/view/ar-metalearn).
@@
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; August 25–29, 2024; Peking, Chinaisbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Meta learningccs: Computing methodologies Generative modeling
@@ 1. Introduction
Literature Challenges: Despite the promising performance of meta-learning methods, several concerns remain in the field. One of them is regarding the task distribution since it closely relates to the model’s generalization evaluation (Yu et al., 2020; Conklin et al., 2021). Overall, task identifiers configure the task, such as the topic type in the corpus for large language models (Brown et al., 2020; Wang et al., 2023b), the amplitude and phase in sinusoid functions, or the degree of freedom in robotic manipulators (Faverjon and Tournassoud, 1987; Anne et al., 2021). However, many studies adopt hand-crafted or simple prior task distributions, such as uniform distributions over task identifiers (Finn et al., 2017; Garnelo et al., 2018a; Rusu et al., 2018). The resulting hand-crafted or naive task samplers might suffer from catastrophic failures in real-world scenarios when distribution shifts occur in the task space.
@@
Proposed Solutions: Rather than exploring fast adaptation strategies, we turn to explicitly create task distribution shifts at a certain level and model the robustness issue with a Stackelberg game (Osborne et al., 2004). To this end, we utilize normalizing flows to parameterize the distribution adversary in Fig. (1) for task distribution generation and the meta learner for fast adaptation in the presence of distribution shifts.
@@
Outline & Primary Contributions: The remainder starts with related work in Section (2). We define the notation and recap fundamentals in Section (3). Then, we present the game-theoretical framework to handle constrained task distribution shifts and robustify fast adaptation in Section (4). The quantitative analysis is conducted in Section (5), followed by conclusions and limitations. In primary, our contributions are:
@@ 2. Literature Review
Meta Learning. Meta learning is a learning paradigm that considers a distribution over tasks. The key is to pursue strategies for leveraging past experiences and distilling extracted knowledge into unseen tasks with a few shots of examples (Munkhdalai and Yu, 2017; Hospedales et al., 2021). Currently, there are various families of meta-learning methods. The optimization-based ones, like model agnostic meta-learning (MAML) (Finn et al., 2017) and its extensions (Grant et al., 2018; Rajeswaran et al., 2019; Fallah et al., 2020; Wang et al., 2020), aim at finding a good meta-initialization of model parameters for adapting to all tasks via gradient descent. The deep metrics methods optimize the task representation in a metric space and are superior in few-shot image classification tasks (Snell et al., 2017; Li et al., 2019; Allen et al., 2019; Huang et al., 2021; Yang et al., 2021). Typical context-based methods, e.g., neural processes (NPs) and variants (Garnelo et al., 2018a, b; Wang et al., 2023a; Wang and Van Hoof, 2022; Kim et al., 2018; Gordon et al., 2019; Wang and van Hoof, 2022; Requeima et al., 2019; Shen et al., 2023), constitute the deep latent variable model as the stochastic process to accomplish tasks. Besides, memory-augmented networks (Santoro et al., 2016), hyper-networks (Ha et al., 2016), and so forth are designed for meta-learning purposes.
@@
Robustness in Meta Learning. In most previous work, the task distribution is fixed in the training set-up. In order to robustify the fast adaptation performance, a couple of learning strategies or principles emerge. Increasing the robustness to worst cases is a commonly seen consideration in meta-learning, and these scenarios include input noise, parameter perturbation, and task distributions (Olds, 2015; Kurakin et al., 2016; Liu et al., 2018; Zhang et al., 2020; Tay et al., 2022; Li et al., 2022). To alleviate the effects of adversarial examples in few-shot image classification, Goldblum et al. (2019) meta-train the model in an adversarial way. To handle the distribution mismatch between training and testing tasks, Zhang et al. (2021) adopt the adaptive risk minimization principle to enable fast adaptation. Wang et al. (2023c) propose to optimize the expected tail risk in meta-learning and witness the increase of robustness in proportional worst cases. Ours is a variant of a distributionally robust framework (Wiesemann et al., 2014), and we seek equilibrium for fast adaptation.
@@
Task Distribution Studies in Meta Learning. Task distributions are directly related to the generalization capability of meta-learning models, attracting increasing attention recently. Aiming to alleviate task overfitting, Rajendran et al. (2020); Yao et al. (2021a); Murty et al. (2021); Ni et al. (2021) enriches the task space with augmentation techniques. Task relatedness can improve generalization across tasks, Fifty et al. (2021) devises an efficient strategy to group tasks in multi-task training. In (Liu et al., 2020; Yao et al., 2021b), neural task samplers are developed to schedule the probability of task sampling in the context of few-shot classification. To increase the fidelity of generated tasks, Wu et al. (2022) adopts the task representation model and constructs the up-sampling network for meta-training task augmentation. To reduce the required tasks, (Yao et al., 2021b; Lee et al., 2022) takes the task interpolation strategy and shows that the interpolation strategy outperforms the standard set-up. Distinguished from the above, this work takes more interest in explicitly understanding task identifier structures concerning learning performance and cares about fast adaptation robustness under subpopulation shift constraints.
@@ 3.1. Problem Statement
Here, $\bm{\theta}$ refers to the meta-learning model parameters for meta knowledge and fast adaptation. The risk function depends on specific meta-learning methods. For example, in MAML, the form can be $\mathcal{L}(\mathcal{D}_{\tau}^{Q},\mathcal{D}_{\tau}^{S};\bm{\theta}):=\mathcal{L}(\mathcal{D}_{\tau}^{Q};\bm{\theta}-\nabla_{\bm{\theta}}\mathcal{L}(\mathcal{D}_{\tau}^{S};\bm{\theta}))$ in regression, where the gradient update in the bracket reflects fast adaptation.
@@
Meanwhile, the worst-case optimization for meta-learning in Eq. (3) can be treated as a particular instance of Eq. (2) when $\alpha$ sufficiently approaches $1$.
@@ 4. Task Robust Meta Learning under Distribution Shift Constraints
This section starts with the game-theoretic framework for meta-learning, followed by approximate optimization. Fig. (2) shows a diagram of the constructed Stackelberg game. Then we perform theoretical analysis w.r.t. our approach.
Figure 2. Diagram of Adversarial Robust Meta Learning. The proposed framework consists of two players, the distribution adversary and the meta player, in the game of meta-learning. On the left side of the figure: the distribution adversary seeks to transform the distribution from an initial task distribution, e.g., $\mathcal{N}(0,I_{d})$ or $\mathcal{U}[a,b]$, via the neural network parameterized by $\bm{\phi}$ with the purpose of deteriorating meta player’s fast adaptation performance. On the right side of the figure: the meta player parameterized by $\bm{\theta}$ attempts to learn robust strategies for fast adaptation in sampled worst-case tasks (MAML algorithm (Finn et al., 2017) as an illustration).
@@ 4.1. Generate Task Distribution within A Game-Theoretic Framework
As previously mentioned, the role of the distribution adversary attempts to transform the initial task distribution into one that raises challenging task proposals with higher probability. Such a setup drives the evolution of task distributions via adaptively shifting task sampling chance under constraints, which can be more crucial for generalization across risky scenarios. The term $D_{KL}\Big{[}p_{0}(\tau)\parallel p_{\bm{\phi}}(\tau)\Big{]}$ inside Eq. (5) works as regularization to avoid the mode collapse in the generative task distribution. In Fig. (2), the goal of the meta learner retains that of traditional meta-learning, while the distribution adversary continually generates the task distribution shifts along optimization processes.
@@
Here, we take two typical methods, e.g., MAML (Finn et al., 2017) and CNP (Garnelo et al., 2018a), to illustrate the meta learner within the adversarially task robust framework, see Examples (1)/(2) for details.
@@
As the normalizing flow function is invertible, the Definition (3) is to describe the Lipschitz continuity in bi-directions.
@@ 4.2. Solution Concept & Explanations
Further, we can define the resulting equilibrium as a local minimax point (Jin et al., 2020) in adversarially task robust meta-learning, due to the non-convex optimization practice.
@@
Moreover, there exists a clearer interpretation w.r.t. the sequential optimization process and the equilibrium in the Definition (4). The meta learner as the leader first optimizes its parameter $\bm{\theta}$. Then the distribution adversary as the follower updates the parameter $\bm{\phi}$ and explicitly generates the task distribution proposal to challenge adaptation performance. In other words, we expect that meta learners can benefit from generative task distribution shifts regarding the adaptation robustness.
@@ 4.3. Strategies for Finding Equilibrium
The form of stochastic gradients w.r.t. the meta player parameter $\bm{\theta}$ is the meta-learning algorithm specific or model dependent. We refer the reader to Algorithm (1)/(2) as examples.
@@
where the particle $\tau_{k}\sim p_{\bm{\phi}}(\tau)$ denotes the task sampled from the generative task distribution, and $\tau_{k}^{-M}$ means the particle sampled from the initial task distributoin to enable $\texttt{NN}_{\bm{\phi}}(\tau_{k}^{-M})=\tau_{k}$.
@@
Particularly, since the normalizing flow works as the distribution transformer in this work, please refer to Eq. (10) to obtain the derivative of the log-likelihood of the transformed task $\ln p_{\bm{\phi}}(\tau)$ w.r.t. $\bm{\phi}$ inside Eq. (17). For easier analysis, we characterize the iteration sequence in optimization as $\begin{bmatrix}\bm{\theta}_{0}\\ \bm{\phi}_{0}\end{bmatrix}\mapsto\cdots\mapsto\begin{bmatrix}\bm{\theta}_{t}\\ \bm{\phi}_{t}\end{bmatrix}\mapsto\begin{bmatrix}\bm{\theta}_{t+1}\\ \bm{\phi}_{t+1}\end{bmatrix}\mapsto\cdots.$
@@ 4.4. Theoretical Analysis
Built on the deduction of the local Stackelberg equilibrium’s existence and the Remark (2), we further perform analysis on the considered equilibrium $\begin{bmatrix}\bm{\theta}_{*}\\ \bm{\phi}_{*}\end{bmatrix}$, in terms of learning dynamics using the alternating GDA. For notation simplicity, we denote the block terms inside the Hessian matrix $\mathbf{H}_{*}:=\nabla^{2}\mathcal{J}(\bm{\theta}_{*},\bm{\phi}_{*})$ around $[\bm{\theta}_{*},\bm{\phi}_{*}]^{T}$ as $\begin{bmatrix}\nabla^{2}_{\bm{\theta}\bm{\theta}}\mathcal{J}&\nabla^{2}_{\bm{\theta}\bm{\phi}}\mathcal{J}\\ \nabla^{2}_{\bm{\phi}\bm{\theta}}\mathcal{J}&\nabla^{2}_{\bm{\phi}\bm{\phi}}\mathcal{J}\end{bmatrix}\Big{|}_{[\bm{\theta}_{*},\bm{\phi}_{*}]^{T}}:=\begin{bmatrix}\mathbf{A}&\mathbf{B}\\ \mathbf{B}^{T}&\mathbf{C}\end{bmatrix}$.
@@
Suppose that the Assumption (1) and the function condition of the (local) Stackelberg equilibrium $\Delta(\mathbf{A},\mathbf{B},\mathbf{C},\gamma_{1},\gamma_{2})<\frac{1}{2}$ are satisfied, where norms of the corresponding matrix are involved. Then the following statements hold:
@@
The Theorem (5) clarifies learning rates $\gamma_{1}$ and $\gamma_{2}$’s influence on convergence and the required second-order derivative conditions of the resulting stationary point $[\bm{\theta}_{*},\bm{\phi}_{*}]^{T}$. And when the game arrives at convergence, the local Stackelberg equilibrium is the best response to these two players, which is at least a local min-max solution to Eq. (5).
@@
We refer the reader to Appendix (F) for formal Theorem (6) and proofs. It reveals the connection between the bound and task complexity $\Upsilon(\mathcal{T})$, and more training tasks from initial distributions decrease the generalization error in adversarially distribution shifts.

### v3

@@ Abstract.
Meta-learning is a practical learning paradigm to transfer skills across tasks from a few examples. Nevertheless, the existence of task distribution shifts tends to weaken meta-learners’ generalization capability, particularly when the training task distribution is naively hand-crafted or based on simple priors that fail to cover critical scenarios sufficiently. Here, we consider explicitly generative modeling task distributions placed over task identifiers and propose robustifying fast adaptation from adversarial training. Our approach, which can be interpreted as a model of a Stackelberg game, not only uncovers the task structure during problem-solving from an explicit generative model but also theoretically increases the adaptation robustness in worst cases. This work has practical implications, particularly in dealing with task distribution shifts in meta-learning, and contributes to theoretical insights in the field. Our method demonstrates its robustness in the presence of task subpopulation shifts and improved performance over SOTA baselines in extensive experiments. The code will be available at the project site (https://sites.google.com/view/ar-metalearn).
@@
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining; August 3–7, 2025; Toronto, ON, Canadaisbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Meta learningccs: Computing methodologies Generative modeling
@@ 1. Introduction
Literature Challenges: Despite the promising adaptation performance in meta-learning, several concerns remain. Among them, the automatically task distribution design is under-explored and challenging in the field, which closely relates to the model’s generalization evaluation (Yu et al., 2020; Conklin et al., 2021).
Overall, task identifiers configure the task, such as the topic type in the corpus for large language models (Brown et al., 2020; Wang et al., 2023b), the amplitude and phase in sinusoid functions, or the degree of freedom in robotic manipulators (Faverjon and Tournassoud, 1987; Anne et al., 2021). Most existing studies adopt simple prior, such as uniform distributions over task identifiers (Finn et al., 2017; Garnelo et al., 2018a; Rusu et al., 2018), or hand-crafted distributions, which heavily rely on domain-specific knowledge difficult to acquire.
Some scenarios even pose more realistic demands for task distributions. In testing an autopilot system, an ideal task distribution deserves more attention on traffic accidents or even generates some while covering typical cases (Tan et al., 2021; Rempe et al., 2022). Similar circumstances also occur during domain randomization for embodied robots (Mehta et al., 2020). These imply that the shift between commonly used task distributions, such as uniform, and the expected testing distributions raises robustness issues and probably causes catastrophic failures when adapting to risk-sensitive scenarios (Mao et al., 2023).
Proposed Solutions: Rather than exploring fast adaptation strategies, we turn to explicitly create task distribution shifts at a certain level and characterize robust fast adaptation with a Stackelberg game (Osborne et al., 2004). To this end, we utilize normalizing flows to parameterize the distribution adversary in Figure 1 for task distribution generation and the meta learner for fast adaptation in the presence of distribution shifts.
@@
Outline & Primary Contributions: The remainder starts with related work in Section 2. We define the notation and recap fundamentals in Section 3. Then, we present the game-theoretical framework to handle constrained task distribution shifts and robustify fast adaptation in Section 4. The quantitative analysis is conducted in Section 5, followed by conclusions and limitations. In primary, our contributions are:
@@ 2. Literature Review
Meta Learning. Meta learning is a learning paradigm that considers a distribution over tasks. The key is to pursue strategies for leveraging past experiences and distilling extracted knowledge into unseen tasks with a few shots of examples (Munkhdalai and Yu, 2017; Hospedales et al., 2021; Chi et al., 2022). Currently, there are various families of meta-learning methods. The optimization-based ones, like model agnostic meta-learning (MAML) (Finn et al., 2017) and its extensions (Grant et al., 2018; Rajeswaran et al., 2019; Fallah et al., 2020; Wang et al., 2020), aim at finding a good meta-initialization of model parameters for adapting to all tasks via gradient descent. The deep metrics methods optimize the task representation in a metric space and are superior in few-shot image classification tasks (Snell et al., 2017; Li et al., 2019; Allen et al., 2019; Huang et al., 2021; Yang et al., 2021). Typical context-based methods, e.g., neural processes (NPs) and variants (Garnelo et al., 2018a; Wang and Van Hoof, 2020; Garnelo et al., 2018b; Wang et al., 2023a; Wang and Van Hoof, 2022; Kim et al., 2018; Gordon et al., 2019; Wang and van Hoof, 2022; Requeima et al., 2019; Shen et al., 2023), constitute the deep latent variable model as the stochastic process to accomplish tasks. Besides, memory-augmented networks (Santoro et al., 2016), hyper-networks (Ha et al., 2016), and so forth are designed for meta-learning purposes.
@@
Robustness in Meta Learning. In most previous work, the task distribution is fixed in the training set-up. In order to robustify the fast adaptation performance, a couple of learning strategies or principles emerge. Increasing the robustness to worst cases is a commonly seen consideration in adaptation, and these scenarios include input noise, parameter perturbation, and task distributions (Olds, 2015; Kurakin et al., 2016; Liu et al., 2018; Zhang et al., 2020; Chi et al., 2021; Tay et al., 2022; Li et al., 2022). To alleviate the effects of adversarial examples in few-shot image classification, Goldblum et al. (2019) meta-train the model in an adversarial way. To handle the distribution mismatch between training and testing tasks, Zhang et al. (2021) adopt the adaptive risk minimization principle to enable fast adaptation. Wang et al. (2023c) propose to optimize the expected tail risk in meta-learning and witness the increase of robustness in proportional worst cases. Ours is a variant of a distributionally robust framework (Wiesemann et al., 2014), and we seek equilibrium for fast adaptation.
@@
Task Distribution Studies in Meta Learning. Task distributions are directly related to the generalization capability of meta-learning models, attracting increasing attention recently. Aiming to alleviate task overfitting, Rajendran et al. (2020); Yao et al. (2021a); Murty et al. (2021); Ni et al. (2021) enrich the task space with augmentation techniques. Task relatedness can improve generalization across tasks, Fifty et al. (2021) devise an efficient strategy to group tasks in multi-task training. In (Liu et al., 2020; Yao et al., 2021b), neural task samplers are developed to schedule the probability of task sampling in the context of few-shot classification. To increase the fidelity of generated tasks, Wu et al. (2022) adopt the task representation model and constructs the up-sampling network for meta-training task augmentation. To reduce the required tasks, (Yao et al., 2021b; Lee et al., 2022) take the task interpolation strategy and shows that the interpolation strategy outperforms the standard set-up. Distinguished from the above, this work takes more interest in explicitly understanding task identifier structures concerning learning performance and cares about fast adaptation robustness under subpopulation shift constraints. Optimizing the task distribution might reserve the potential to improve generative performance in large models (Chi et al., 2024).
@@ 3.1. Problem Statement
Here, $\bm{\theta}$ refers to the meta-learning model parameters for meta knowledge and fast adaptation. The risk function depends on specific meta-learning methods. For example, in MAML, the form can be $\mathcal{L}(\mathcal{D}_{\tau}^{Q},\mathcal{D}_{\tau}^{S};\bm{\theta}):=\mathcal{L}(\mathcal{D}_{\tau}^{Q};\bm{\theta}-\lambda\nabla_{\bm{\theta}}\mathcal{L}(\mathcal{D}_{\tau}^{S};\bm{\theta}))$ in regression, where the gradient update with the learning rate $\lambda$ in the bracket reflects fast adaptation.
@@
It is worth noting that $p_{\alpha}(\tau;\bm{\theta})$ is non-differentiable and $\theta$-dependent with no closed-form. Meanwhile, the worst-case optimization for meta-learning in Eq. (3) can be treated as a particular instance of Eq. (2) when $\alpha$ sufficiently approaches $1$.
@@ 4. Task Robust Meta Learning under Distribution Shift Constraints
This section starts with the game-theoretic framework for meta-learning, followed by approximate optimization. Figure 2 shows a diagram of the constructed Stackelberg game. Then we perform theoretical analysis w.r.t. our approach.
Figure 2. Diagram of Adversarially Task Robust Meta Learning. The proposed framework consists of two players, the distribution adversary and the meta player, in the game of meta-learning. On the left side of the figure: the distribution adversary seeks to transform the distribution from an initial task distribution, e.g., $\mathcal{N}(0,I_{d})$ or $\mathcal{U}[a,b]$, via the neural network parameterized by $\bm{\phi}$ with the purpose of deteriorating meta player’s fast adaptation performance. On the right side of the figure: the meta player parameterized by $\bm{\theta}$ attempts to learn robust strategies for fast adaptation in sampled worst-case tasks (MAML algorithm (Finn et al., 2017) as an illustration).
@@ 4.1. Generate Task Distribution within A Game-Theoretic Framework
As previously mentioned, the role of the distribution adversary attempts to transform the initial task distribution into one that raises challenging task proposals with higher probability. Such a setup drives the evolution of task distributions via adaptively shifting task sampling chance under constraints, which can be more crucial for generalization across risky scenarios. The term $D_{KL}\Big{[}p_{0}(\tau)\parallel p_{\bm{\phi}}(\tau)\Big{]}$ inside Eq. (5) works as regularization to avoid the mode collapse in the generative task distribution. In Figure 2, the goal of the meta learner retains that of traditional meta-learning, while the distribution adversary continually generates the task distribution shifts along optimization processes.
@@
Here, we take two typical methods, e.g., MAML (Finn et al., 2017) and CNP (Garnelo et al., 2018a), to illustrate the meta learner within the adversarially task robust framework, see Examples 1/2 for details.
@@
As the normalizing flow function is invertible, the Definition 3 is to describe the Lipschitz continuity in bi-directions.
@@ 4.2. Solution Concept & Explanations
Moreover, we can define the resulting equilibrium as a local minimax point (Jin et al., 2020) in adversarially task robust meta-learning, due to the non-convex optimization practice.
@@
Moreover, there exists a clearer interpretation w.r.t. the sequential optimization process and the equilibrium in the Definition 4. The meta learner as the leader first optimizes its parameter $\bm{\theta}$. Then the distribution adversary as the follower updates the parameter $\bm{\phi}$ and explicitly generates the task distribution proposal to challenge adaptation performance. In other words, we expect that meta learners can benefit from generative task distribution shifts regarding the adaptation robustness.
@@ 4.3. Strategies for Finding Equilibrium
The form of stochastic gradients w.r.t. the meta player parameter $\bm{\theta}$ is the meta-learning algorithm specific or model dependent. We refer the reader to Algorithm 1/2 as examples.
@@
where the particle $\tau_{k}\sim p_{\bm{\phi}}(\tau)$ denotes the task sampled from the generative task distribution, and $\tau_{k}^{-M}$ means the particle sampled from the initial task distribution to enable $\texttt{NN}_{\bm{\phi}}(\tau_{k}^{-M})=\tau_{k}$.
@@
Particularly, since the normalizing flow works as the distribution transformation in this work, please refer to Eq. (10) to obtain the derivative of the log-likelihood of the transformed task $\ln p_{\bm{\phi}}(\tau)$ w.r.t. $\bm{\phi}$ inside Eq. (17). For easier analysis, we characterize the iteration sequence in optimization as $\begin{bmatrix}\bm{\theta}_{0}\\ \bm{\phi}_{0}\end{bmatrix}\mapsto\cdots\mapsto\begin{bmatrix}\bm{\theta}_{t}\\ \bm{\phi}_{t}\end{bmatrix}\mapsto\begin{bmatrix}\bm{\theta}_{t+1}\\ \bm{\phi}_{t+1}\end{bmatrix}\mapsto\cdots.$
@@ 4.4. Theoretical Analysis
Built on the deduction of the local Stackelberg equilibrium’s existence and the Remark 2, we further perform analysis on the considered equilibrium $\begin{bmatrix}\bm{\theta}_{*}\\ \bm{\phi}_{*}\end{bmatrix}$, in terms of learning dynamics using the alternating GDA. For notation simplicity, we denote the block terms inside the Hessian matrix $\mathbf{H}_{*}:=\nabla^{2}\mathcal{J}(\bm{\theta}_{*},\bm{\phi}_{*})$ around $[\bm{\theta}_{*},\bm{\phi}_{*}]^{T}$ as $\begin{bmatrix}\nabla^{2}_{\bm{\theta}\bm{\theta}}\mathcal{J}&\nabla^{2}_{\bm{\theta}\bm{\phi}}\mathcal{J}\\ \nabla^{2}_{\bm{\phi}\bm{\theta}}\mathcal{J}&\nabla^{2}_{\bm{\phi}\bm{\phi}}\mathcal{J}\end{bmatrix}\Big{|}_{[\bm{\theta}_{*},\bm{\phi}_{*}]^{T}}:=\begin{bmatrix}\mathbf{A}&\mathbf{B}\\ \mathbf{B}^{T}&\mathbf{C}\end{bmatrix}$.
@@
Suppose that the Assumption 1 and the function condition of the (local) Stackelberg equilibrium $\Delta(\mathbf{A},\mathbf{B},\mathbf{C},\gamma_{1},\gamma_{2})<\frac{1}{2}$ are satisfied, where norms of the corresponding matrix are involved. Then the following statements hold:
@@
The Theorem 5 clarifies learning rates $\gamma_{1}$ and $\gamma_{2}$’s influence on convergence and the required second-order derivative conditions of the resulting stationary point $[\bm{\theta}_{*},\bm{\phi}_{*}]^{T}$. And when the game arrives at convergence, the local Stackelberg equilibrium is the best response to these two players, which is at least a local min-max solution to Eq. (5).
@@
We refer the reader to Appendix F for formal Theorem 6 and proofs. It reveals the connection between the bound and task complexity $\Upsilon(\mathcal{T})$, and more training tasks from initial distributions decrease the generalization error in adversarially distribution shifts.